OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6 ???? i dont know how to do this at all

sam (.sam.):

Where you've stucked?

OpenStudy (anonymous):

i distributed, but i dont know what to do with the +6.

OpenStudy (kymber):

Can you show what you've done thus far?

Parth (parthkohli):

1) Distribute first. \(\Large \color{MidnightBlue}{\Rightarrow 3x + 3 = -2x + 2 + 6 }\) 2)Operate on like terms. \(\Large \color{MidnightBlue}{\Rightarrow 3x + 3 = -2x + 8 }\) 3) Add 2x to both sides. \(\Large \color{MidnightBlue}{\Rightarrow 5x + 3 = 8 }\) 4) Subtract 3 from both sides. \(\Large \color{MidnightBlue}{\Rightarrow 5x = 5 }\) 5) Divide both sides by 5. \(\Large \color{MidnightBlue}{\Rightarrow x = 1 }\)

sam (.sam.):

Multiply -2 by each term inside the parentheses. 3(x+1)=-2x+2+6 Add 6 to 2 to get 8. 3(x+1)=-2x+8 3x+3=-2x+8 solve from here?

Parth (parthkohli):

You must isolate the x.

OpenStudy (anonymous):

yes thank you very much (: